OpenStudy (anonymous):

Anyone tryna help me with 5 questions for a medal ?

pooja195 (pooja195):

Welcome to openstudy :) Can you post your question?

OpenStudy (anonymous):

Im not new but thank you The table shows the ratio of caramel corn to cheddar corn in a bag. What is the constant of proportionality? Cheddar Corn Caramel Corn 10 15 20 30 30 45

OpenStudy (anonymous):

\[\frac{15}{10}=\frac{30}{20}=\frac{45}{30}=?\]

OpenStudy (anonymous):

Im not good with fractions I need a minute to figure it out

OpenStudy (anonymous):

50 over 35 ?

OpenStudy (anonymous):

the idea is to take any one of those equivalent fractions and reduce it to lowest terms

OpenStudy (anonymous):

Oh 5 over 10 ?

OpenStudy (anonymous):

no reduce \[\frac{15}{10}\] by dividing top and bottom by the common factor of 5

OpenStudy (anonymous):

3 over 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or \(\frac{3}{2}\) or \(1\tfrac{1}{2}\) or even \(1.5\)

OpenStudy (anonymous):

Could you help with some more (thank you)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

best to post in a new thread, hard to keep scrolling down

OpenStudy (anonymous):

Okay